# VBench-2.0: Advancing Video Generation Benchmark Suite for Intrinsic Faithfulness

### v1

@@ Abstract
Video generation has advanced significantly, evolving from producing unrealistic outputs to generating videos that appear visually convincing and temporally coherent. 00footnotetext: equal contributions. 🖂corresponding authors.   Code is available To evaluate these video generative models, benchmarks such as VBench have been developed to assess their faithfulness, measuring factors like per-frame aesthetics, temporal consistency, and basic prompt adherence. However, these aspects mainly represent superficial faithfulness, which focus on whether the video appears visually convincing rather than whether it adheres to real-world principles. While recent models perform increasingly well on these metrics, they still struggle to generate videos that are not just visually plausible but fundamentally realistic. To achieve real “world models” through video generation, the next frontier lies in intrinsic faithfulness to ensure that generated videos adhere to physical laws, commonsense reasoning, anatomical correctness, and compositional integrity. Achieving this level of realism is essential for applications such as AI-assisted filmmaking and simulated world modeling.
@@
1 Introduction
Figure 2: VBench-2.0 Evaluation Results of SOTA Models. The figure presents the evaluation results of four recent state-of-the-art video generation models across 18 VBench-2.0 dimensions. The results are normalized per dimension for a clearer comparison. For detailed numerical results, refer to Table 2.
@@ 1 Introduction
Our evaluation provides comprehensive insights into the strengths and weaknesses of state-of-the-art video generative models. While recent models demonstrate emerging abilities in human anatomy, consistency, and some degree of novel creativity, they still struggle with generating complex plots, handling simple dynamic changes in objects, and remain unstable in commonsense reasoning, highlighting key open challenges in video generation towards synthesizing the world with intrinsic faithfulness. We provide an in-depth discussion on possible causes, potential solutions, and inherent trade-offs in Section 5.
@@
2 Related Works
@@
Video Generative Models. With the advancements in diffusion models [63, 65, 27, 64, 90, 2, 17, 50, 31, 13, 14, 28], variational autoencoder-based compression techniques [37, 77, 16, 56, 87], and transformer architectures [15, 54], video generation has emerged as one of the most dynamic frontiers in artificial intelligence research. Prior to Sora’s breakthrough, predominant text-to-video models primarily focused on synthesizing short video clips (2-3 seconds duration) [5, 80, 23, 29, 48, 24, 95, 88, 21, 3, 78, 36, 7, 86, 57, 53, 68] through incremental improvements in visual fidelity and temporal consistency. Sora [53] pioneered the scaling paradigm in video generation by demonstrating unprecedented model capacity through large-scale training, paving the way for the development of next-generation video foundation models [86, 59, 69, 72, 73, 71, 1, 70, 18, 61, 55] that achieve remarkable visual quality and robust spatiotemporal coherence. They have shifted focus toward enhancing video generation adhere to deeper principles such as physical laws and commonsense reasoning that focuses more on action continuity and realistic physical-world perception [86, 53, 1], or high-quality human-centric generation with creativity potential [69, 72]. Existing benchmarks cannot systematically evaluate these new explorations, and VBench-2.0 takes the initiative to provide a comprehensive benchmark for evaluating emerging capabilities towards the goal of achieving intrinsic faithfulness through video generation.
@@
Evaluation of Video Generative Models. Initially, video generative models primarily relied on conventional evaluation metrics such as Fréchet inception distance (FID) [26], Inception Score (IS) [60], and Fréchet video distance (FVD) [76]. However, these metrics provided limited insight into the diverse and complex capabilities of modern video generation. Recent evaluation frameworks [32, 33, 46, 45, 89] such as VBench [32, 33] introduced a more structured approach by disentangling evaluation into multiple capability dimensions, enabling more detailed and interpretable assessments. These benchmarks focus on fundamental technical attributes such as per-frame quality, temporal consistency, and basic prompt adherence. However, as models continue to improve, certain dimensions within VBench begin to saturate, necessitating broader evaluation scope that assess deeper aspects of intrinsic faithfulness in video generation. To address this, specialized benchmarks have emerged. PhyGenBench [49] evaluates a model’s understanding of physical laws through Vision-Language Models (VLMs). T2V-CompBench [66] assesses compositionality, including motion, actions, spatial relationships, and attributes. StoryEval [81] focuses on storytelling capabilities by aggregating the responses from two VLMs. Unlike prior benchmarks, which either focus on fundamental capabilities [46, 32, 33] or specific emerging domains [49, 66], VBench-2.0 introduces a comprehensive framework to systematically evaluate next-generation video generation capabilities, bridging the gap in the evolving landscape of video generation.
@@
Table 1: Comparison of Video Generation Benchmarks. We compare existing video generation benchmarks based on their evaluation aspects. VBench-2.0 is the first comprehensive benchmark to assess intrinsic faithfulness in video generation, complementing VBench [32, 33]. Detailed aspects include per-frame quality (Frame Wise), temporal consistency (Temp Cons), adherence to simple prompts (Simp Pmpt), compositional creativity (Comp Crea), commonsense reasoning (Com Sense), physics-based realism (Phy), human anatomy (Human Anat), and adherence to complex prompts (Cplx Pmpt).
@@
[…]
Pmpt
@@
3 VBench-2.0 Suite for Intrinsic Faithfulness
In this section, we introduce the evaluation framework of VBench-2.0. Section 3.1 presents the five key evaluation dimensions and their respective assessment methods. Unlike VBench, which primarily evaluates superficial faithfulness, VBench-2.0 introduces a suite of tests to assess intrinsic faithfulness, focusing on deeper properties such as physics, commonsense, and creativity, human, and controllability. To ensure robust evaluation, we integrate multiple assessment methodologies, including LLM-assisted text alignment, video-based multi-question answering, and specialist models trained for anomaly detection. Our prompt suite is introduced in Section 3.2.
3.1 Evaluation Dimension Suite
@@
where V is the generated video, T denotes the text prompt or crafted reference, and $S_{v}$ and $S_{l}$ are the system prompts for the VLM and LLM, respectively. The LLM outputs a binary judgment (“yes” or “no”) that is discretized to a score of 1 or 0 to reflect caption-reference matching. This scheme excels in semantic understanding dimensions like Complex Plot and Human Interaction, where VLMs usually struggle with high-level interpretation, but LLMs demonstrate stronger reasoning capabilities. Thus, we decouple caption generation (by VLMs) from semantic alignment (by LLMs) to improve evaluation reliability.
@@
where Q is a set of multiple questions. For example, in the Dynamic Attribute dimension focusing on color changes, we may ask: 1. Initially, is the color of the river mostly blue? 2. Finally, is the color of the river mostly brown? 3. Does the color of the river change? The answer to each question is binary (“yes” or “no”), and scores are either averaged or awarded only if all responses are correct, depending on the dimension’s scoring scheme. This scheme is particularly effective for surface-level visual understanding, where modern VLMs can confidently answer targeted queries without requiring high-level semantic reasoning.
3.1.1 Human Fidelity
@@
We evaluate both the structural correctness and temporal consistency of human figures in generated videos. Structural issues commonly seen in current video generation models, such as sudden turns or the “thousand-hand yoga” effect, are considered. For temporal consistency, we assess the consistency human identity and clothing across frames.
Human Anatomy. We assess structural realism by identifying unnatural per-frame deformations in hands, faces, and bodies. To achieve this, we train three anomaly detection models using a pre-trained ViT-based backbone [82], using a curated dataset of 150k real and generated human frames.
Human Temporal Consistency - Clothes. Clothing consistency is assessed using video-based multi-question answering to ensure outfits remain stable throughout video.
Human Temporal Consistency - Identity. Identity consistency is evaluated by measuring facial feature similarity using ArcFace [11], with face detection performed by RetinaFace [12].
3.1.2 Creativity
@@
We evaluate creativity by analyzing a model’s ability to generate diverse outputs and complex compositions beyond real-world constraints.
@@
Diversity. Given a text prompt, we sample 20 videos from a model, and measure inter-sample variation using style and content diversity metrics, computed from pre-trained VGG-19 [62] feature representations.
@@
Composition. We assess species combination, single-entity actions, and multi-entity interactions using a structured video-based multi-question answering pipeline. This approach measures whether the model can generate novel and uncommon compositions.
@@
3.1.3 Controllability
We evaluate a model’s ability to follow complex prompts and simulate dynamic changes during video generation. This dimension measures how accurately the model can render specific entities, events, content, and camera movements in response to detailed textual instructions.
Entity - Dynamic Spatial Relationship. We assess whether models accurately reposition objects in response to spatial instructions (e.g., “A dog is on the left of a sofa, then the dog runs to the front of the sofa.”) using video-based multi-question answering.
Entity - Dynamic Attribute. We test whether models can modify attributes (e.g., color, size, texture) mid-video using the video-based multi-question answering pipeline.
Event - Motion Order Understanding. We evaluate whether models generate several actions or motions in the specified order. We use the text description alignment pipeline to measure whether the generated motion sequence matches the text prompt.
Event - Human Interaction. We assess whether two humans can interact (e.g., “One person hands an object to another”) based on the text prompt. We use the text description alignment pipeline.
Content - Complex Landscape. We evaluate whether models faithfully follow long-form landscape descriptions (150+ words) that include multiple scene transitions driven by camera movements. We assess adherence using text description alignment with landscape-specific system prompt.
Content - Complex Plot. We test a model’s ability to construct multi-scene narratives from prompts describing multi-stage events (e.g., a four-act story with 150+ words). We evaluate plot consistency using text description alignment, similar to the landscape evaluation.
Camera Motion. We evaluate whether models can generate specified camera movements. We extend VBench++ [33]’s camera motion taxonomy to nine types, adding “Orbit” and “Oblique shot, airborne dolly movement.” The generated camera motion is assessed via point tracking with CoTracker-v2 [34] and carefully tailored heuristics.

### v2

@@ Abstract
Video generation has advanced significantly, evolving from producing unrealistic outputs to generating videos that appear visually convincing and temporally coherent. To evaluate these video generative models, benchmarks such as VBench have been developed to assess their faithfulness, measuring factors like per-frame aesthetics, temporal consistency, and basic prompt adherence. However, these aspects mainly represent superficial faithfulness, which focus on whether the video appears visually convincing rather than whether it adheres to real-world principles. While recent models perform increasingly well on these metrics, they still struggle to generate videos that are not just visually plausible but fundamentally realistic. To achieve real “world models” through video generation, the next frontier lies in intrinsic faithfulness to ensure that generated videos adhere to physical laws, commonsense reasoning, anatomical correctness, and compositional integrity. Achieving this level of realism is essential for applications such as AI-assisted filmmaking and simulated world modeling.
@@
Index Terms:
Video Generative Models, Evaluation Benchmark.
@@
I Introduction
@@
Figure 2: VBench-2.0 Evaluation Results of SOTA Models. The figure presents the evaluation results of four recent state-of-the-art video generation models across 18 VBench-2.0 dimensions. The results are normalized per dimension for a clearer comparison. For detailed numerical results, refer to Table II.
@@
Our evaluation provides comprehensive insights into the strengths and weaknesses of state-of-the-art video generative models. While recent models demonstrate emerging abilities in human anatomy, consistency, and some degree of novel creativity, they still struggle with generating complex plots, handling simple dynamic changes in objects, and remain unstable in commonsense reasoning, highlighting key open challenges in video generation towards synthesizing the world with intrinsic faithfulness. We provide an in-depth discussion on possible causes, potential solutions, inherent trade-offs, and future work in Section V.
@@
II Related Works
II-A Video Generative Models
@@
With the advancements in diffusion models [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], variational autoencoder-based compression techniques [45, 46, 47, 48, 49], and transformer architectures [50, 51], video generation has emerged as one of the most dynamic frontiers in artificial intelligence research. Prior to Sora’s breakthrough, predominant text-to-video models primarily focused on synthesizing short video clips (2-3 seconds duration) [2, 21, 18, 24, 52, 53, 54, 55, 56, 57, 22, 58, 23, 59, 60, 28, 32] through incremental improvements in visual fidelity and temporal consistency. Sora [28] pioneered the scaling paradigm in video generation by demonstrating unprecedented model capacity through large-scale training, paving the way for the development of next-generation video foundation models [59, 30, 29, 31, 61, 62, 20, 63, 64, 65, 66] that achieve remarkable visual quality and robust spatiotemporal coherence. They have shifted focus toward enhancing video generation adhere to deeper principles such as physical laws and commonsense reasoning that focuses more on action continuity and realistic physical-world perception [59, 28, 20], or high-quality human-centric generation with creativity potential [29, 31]. Existing benchmarks cannot systematically evaluate these new explorations, and VBench-2.0 takes the initiative to provide a comprehensive benchmark for evaluating emerging capabilities towards the goal of achieving intrinsic faithfulness through video generation.
II-B Evaluation of Video Generative Models
@@
Initially, video generative models primarily relied on conventional evaluation metrics such as Fréchet inception distance (FID) [67], Inception Score (IS) [68], and Fréchet video distance (FVD) [69]. However, these metrics provided limited insight into the diverse and complex capabilities of modern video generation. Recent evaluation frameworks [25, 26, 27, 70, 71] such as VBench [25, 26] introduced a more structured approach by disentangling evaluation into multiple capability dimensions, enabling more detailed and interpretable assessments. These benchmarks focus on fundamental technical attributes such as per-frame quality, temporal consistency, and basic prompt adherence. However, as models continue to improve, certain dimensions within VBench begin to saturate, necessitating broader evaluation scope that assess deeper aspects of intrinsic faithfulness in video generation. To address this, specialized benchmarks have emerged. PhyGenBench [72] evaluates a model’s understanding of physical laws through Vision-Language Models (VLMs). T2V-CompBench [73] assesses compositionality, including motion, actions, spatial relationships, and attributes. StoryEval [74] focuses on storytelling capabilities by aggregating the responses from two VLMs. Unlike prior benchmarks, which either focus on fundamental capabilities [27, 25, 26] or specific emerging domains [72, 73], VBench-2.0 introduces a comprehensive framework to systematically evaluate next-generation video generation capabilities, bridging the gap in the evolving landscape of video generation.
@@
TABLE I: Comparison of Video Generation Benchmarks. We compare existing video generation benchmarks based on their evaluation aspects. VBench-2.0 is the first comprehensive benchmark to assess intrinsic faithfulness in video generation, complementing VBench [25, 26]. Detailed aspects include per-frame quality (Frame Wise), temporal consistency (Temp Cons), adherence to simple prompts (Simp Pmpt), compositional creativity (Comp Crea), commonsense reasoning (Com Sense), physics-based realism (Phy), human anatomy (Human Anat), and adherence to complex prompts (Cplx Pmpt).
@@
Frame Wise Temp Cons Simp Pmpt Comp Crea Com Sense Phy Human Anat Cplx Pmpt
@@
III VBench-2.0 Suite for Intrinsic Faithfulness
In this section, we introduce the evaluation framework of VBench-2.0. Section III-A presents the five key evaluation dimensions and their respective assessment methods. Unlike VBench, which primarily evaluates superficial faithfulness, VBench-2.0 introduces a suite of tests to assess intrinsic faithfulness, focusing on deeper properties such as physics, commonsense, and creativity, human, and controllability. To ensure robust evaluation, we integrate multiple assessment methodologies, including LLM-assisted text alignment, video-based multi-question answering, and specialist models trained for anomaly detection. Prompt suite is introduced in Section III-B and Section III-C describes the human annotation pipeline of VBench-2.0, including the collection and processing of video data, labeling of human annotations, and the definition of scoring criteria.
III-A Evaluation Dimension Suite
@@
where V is the generated video, T denotes the text prompt or crafted reference, and $S_{v}$ and $S_{l}$ are the system prompts for the VLM and LLM, respectively. The LLM outputs a binary judgment (“yes” or “no”) that is discretized to a score of 1 or 0 to reflect caption-reference matching. This scheme excels in semantic understanding dimensions like Complex Plot and Human Interaction, where VLMs usually struggle with high-level interpretation, but LLMs demonstrate stronger reasoning capabilities. Thus, we decouple caption generation (by VLMs) from semantic alignment (by LLMs) to improve evaluation reliability. Unless otherwise specified, we adopt LLaVA-Video-7B [75] as our VLM model and Qwen2.5-7B-Instruct [76] as LLM model.
@@
where Q is a set of multiple questions. For example, in the Dynamic Attribute dimension focusing on color changes, we may ask: 1. Initially, is the color of the river mostly blue? 2. Finally, is the color of the river mostly brown? 3. Does the color of the river change? The answer to each question is binary (“yes” or “no”), and scores are either averaged or awarded only if all responses are correct, depending on the dimension’s scoring scheme. This scheme is particularly effective for surface-level visual understanding, where modern VLMs can confidently answer targeted queries without requiring high-level semantic reasoning. Unless otherwise specified, we adopt LLaVA-Video-7B [75] as our VLM model.
Figure 3: Human Anatomy Detector Framework. The [cls] token is used to aggregate information related to anomalies in the image. The human body, hand, and face detectors share the same pipeline with different data. The input is frame level and pre-detected.
@@
1) Human Fidelity: We evaluate both the structural correctness and temporal consistency of human figures in generated videos. Structural issues commonly seen in current video generation models, such as sudden turns or the “thousand-hand yoga” effect, are considered. For temporal consistency, we assess the consistency of human identity and clothing across frames.
Figure 4: Visualization for Human Anatomy.
(1-a) Human Anatomy. This dimension focuses on detecting potential anomalies in the appearance and structure of the human body in videos. We use a pre-trained ViT-base model [77] to train three anomaly detection models, which will detect the human in each frame and judge the anomaly score of each human, targeting the human body, hands, and faces, respectively. The classification token (cls_token) is passed through an MLP to produce binary outputs. We show the pipeline in Figure 3 and the dataset consists of two components:
  • Real samples: We collected approximately 1,000 real motion videos from the web and used the YOLO-World [78] detection model to extract pre-frame patches of human bodies, hands, and faces as positive samples.
  • Generated samples: We generated around 1,000 videos related to human motion using CogVideo [24, 59] and HunyuanVideo [31]. These were processed with YOLO-World to extract image patches, followed by meticulous manual annotations for training. Additionally, negative samples were sampled from the HumanRefiner [79] dataset, focusing on human body, hands, and face.
As a result, we collect 150k labeled real and generated human frame-level data. During inference, we first applied the YOLO-World model to detect all human instances in the input video. For each detected human region, we further detected hands and faces, extracting corresponding patches as inputs. A human instance is flagged as abnormal if any of the three models consider this human as an anomaly and the final score is the percentage of frames that are not flagged as abnormal.
(1-b) Human Temporal Consistency - Clothes. Clothing consistency is assessed using video-based multi-question answering to assess whether outfits remain stable throughout video. Note that we do not use traditional methods such as feature similarity calculation to handle this problem based on two considerations: 1) Current models often generate characters with their clothing partially obscured by objects or with changes in visible body parts (e.g., the upper body transforming into the lower body). In such cases, traditional algorithms are unable to address these cross-temporal judgment issues. 2) LLaVA-video-7B has strong clothing color perception and a certain degree of cross-temporal memory capability. Therefore, we use it to evaluate this dimension.
(1-c) Human Temporal Consistency - Identity. Identity consistency is evaluated by measuring facial feature similarity using ArcFace [80], with face detection performed by RetinaFace [81]. Specifically, we select the first frame of the video as the anchor by default, and all subsequent frames are compared to the first frame to calculate similarity. However, time periods where there are multiple people or no one present are not taken into consideration. Note that after a scene change, whether it is still the same person is taken into consideration, which represents a more challenging scenario.
@@
2) Creativity: We evaluate creativity by analyzing a model’s ability to generate diverse outputs and complex compositions beyond real-world constraints.
Figure 5: Visualization for Diversity.
@@
(2-a) Diversity. Given a text prompt, we sample 20 videos from a model, and measure inter-sample variation using style and content diversity metrics, computed from pre-trained VGG-19 [82] feature representations ( i.e., the metric is modified from [83]).
@@
(2-b) Composition. This dimension measures whether the model can generate novel and uncommon compositions and we assess it through three sub-dimensions: species combination, single-entity actions, and multi-entity interactions, using video-based multi-question answering pipeline. We additionally observe that pure VQA can not ensure robust results as current video generation models will generate separate species for the given prompt and VQA is a discrete process that could not consider the separate situations. So in this paper, we omit the cases that show more than one creature in the video by a pre-VQA question “Is there only one creature in the video?”.
3) Controllability: We evaluate a model’s ability to follow complex prompts and simulate dynamic changes during video generation. This dimension measures how accurately the model can render specific entities, events, content, and camera movements in response to detailed textual instructions.
(3-a) Entity - Dynamic Attribute. We test whether models can modify attributes (e.g., color, size, texture) mid-video using the video-based multi-question answering pipeline. We take an example of questions here: “Initially, is the color of leaves mostly red?”; “Finally, is the color of leaves mostly green?”; “Does the color of leaves change?”
(3-b) Entity - Dynamic Spatial Relationship. We assess whether models accurately reposition objects in response to spatial instructions (e.g., “A dog is on the left of a sofa, then the dog runs to the front of the sofa.”) using video-based multi-question answering. The question template is similar to Dynamic Attribute.
(3-c) Event - Motion Order Understanding. We evaluate whether models generate several actions or motions in the specified order. We use the text description alignment pipeline to measure whether the generated motion sequence matches the text prompt. Specifically, we first split the text prompt into two descriptions with action only in turn, then we obtain the video caption by VLM with an additional system prompt “Return the action order in video. Here is the template: ‘1. ; 2. .”’. Then we assess whether both actions match the ground truth descriptions in turn; only when both actions match are they considered correct. This approach cleverly avoids the hallucination problem of LLM in order judgment.
(3-d) Event - Human Interaction. We assess whether two humans can interact (e.g., “One person hands an object to another”) based on the text prompt using text description alignment. The system prompt for VLM is “Describe the human interaction in the video, following the template as [a person xx to another person.]”. After converting the description into a standardized format, the LLM will determine whether the standardized description matches the ground truth. However, the wrong number of people may lead to interaction hallucinations in VLMs, such as treating situations with only one person as interactions between two individuals, so we pre-filter the video with wrong human number with another text description alignment pipeline. Specifically, we obtain the dense description of the video with system prompt “Describe the video in detail” and ask the LLM to determine whether the number of people mentioned in the video description is exactly one.
(3-e) Content - Complex Plot. We test a model’s ability to construct multi-scene narratives from prompts describing multi-stage events (e.g., a five-act story with 150+ words) to meet the demands of future movie-level video generation. We utilize text description alignment to evaluate it. Specifically, we first manually split and summarize the long text prompt into 5 plot descriptions as ground truth. Then we feed the video into VLM to obtain 5 video captions with the system prompt “Return the plot in video. Here is the template: [1. ; 2. ; 3. ; 4. ; 5. .]” ( i.e., note that when dealing with long videos at the minute level, VLMs may make errors in their point-by-point summarization. Therefore, we additionally employ an LLM for post-processing verification to correct captions with incorrect numbering). Finally, we sequentially match the corresponding text and captions in the order of the plot. If the LLM determines that a plot element appears in the caption, we proceed to evaluate the next plot element; otherwise, the evaluation stops. We discretize the plot elements to calculate the accurate score, simulating the effect of LLM scoring.
Figure 6: Visualization for Complex Landscape.
(3-f) Content - Complex Landscape. We evaluate whether models faithfully follow long-form landscape descriptions (150+ words) that include multiple scene transitions driven by camera movements. We assess adherence using text description alignment with landscape-specific system prompt, similar to Complex Plot.
(3-g) Camera Motion. We evaluate whether models can generate specified camera movements. We extend VBench++ [26]’s camera motion taxonomy to nine types, adding “Orbit” and “Oblique shot, airborne dolly movement.” The generated camera motion is assessed via point tracking with CoTracker-v2 [84] and carefully tailored heuristics (e.g., different camera motions exhibit distinct behaviors at different points. For example, pan left results in all points shifting to the right.).
4) Physics: The Physics dimension evaluates whether video generation models adhere to real-world physical principles. We assess two key areas: 1) State Change, which examines how well models simulate mechanical, thermal, and material transformations, and 2) Geometry, which evaluates the 3D consistency of objects and scenes across different frames. We extend prior benchmarks (e.g., PhyGenBench [72]) by increasing physics scenario difficulty and significantly improving evaluation accuracy with tailored pipelines.
Figure 7: Visualization for Mechanics.
(4-a) State Change - Mechanics. We evaluate whether models follow basic mechanical principles such as gravity, buoyancy, and stress. This is done using video-based multi-question answering. Unlike PhyGenBench, which relies on abstract physics concepts, we prompt GPT-4o to generate explicit visual descriptions of expected physical behavior (e.g., unlike PhyGenBench that uses terminologies to describe the physical phenomena “The liquid’s behavior aligns with the microgravity environment, floating freely, and forming natural blobs without noticeable distortion.”, we describe the physical phenomena’s visual results “The liquid floating, and forming blobs”)). Notably, the GPT-4o-generated descriptions are based solely on the text prompt, remain fixed during evaluation, and do not compromise evaluation reproducibility. To ensure a focused assessment of state changes, we also apply a pre-filtering step to exclude cases where the initial state of the generated video does not align with the prompt.
(4-b) State Change - Thermotics. We evaluate how well the models simulate state transitions such as vaporization, liquefaction, and sublimation. To increase complexity, we introduce temperature-specific prompts (e.g., “A timelapse captures dry ice transforming at -90°C”). Evaluation follows the same video-based multi-question answering approach used in the Mechanics dimension.
(4-c) State Change - Material. We evaluate whether the models correctly depict color mixing, hardness, combustion, and solubility. Similar to Mechanics and Thermotics, we use video-based multi-question answering, testing whether generated videos properly follow material properties.
(4-d) Geometry - Multi-View Consistency. Geometry is a critical aspect for 3D/4D video generation, ensuring that generated entities and scenes maintain structural consistency when viewed from different angles or as the camera moves. However, we do not have access to explicit 3D ground truth of the generated videos, making direct 3D validation infeasible. Instead, we assess multi-view consistency using two complementary metrics: 1) Feature Matching Stability – Measures how well objects retain their geometric consistency across frames, inspired by [85], and 2) Camera Motion Speed – Accounts for the effect of motion strengths on feature stability, ensuring fair cross-model comparisons. Specifically, we follow [85] to extract frame-level keypoint features using SIFT [86], efficiently match points across frames with FLANN [87], and eliminate incorrect matches using RANSAC [88]. We then estimate camera motion speed using RAFT [89], and adjust feature-matching frame intervals based on the motion strength and the video frame rate ( i.e., FPS and camera motion speed will influence the score of feature matching).
5) Commonsense: We assess commonsense reasoning in video generation models across two key aspects: 1) Motion Rationality, evaluating whether generated motions are physically plausible and correctly executed, and 2) Instance Preservation, ensuring that abnormal entity states (entity sudden merging, splitting, appearing, and disappearing) do not exist.
Figure 8: Visualization for Instance Preservation.
@@
(5-a) Motion Rationality. We evaluate whether the generated motion leads to the correct real-world consequences. A prevalent issue in video generation is the presence of fake motions that visually appear correct but lack expected effects on the environment. For instance: 1) Fake eating: A person bites into food, but the food remains unchanged. 2) Fake walking: A character moves their legs, but does not actually progress forward. and 3) Fake cutting: A knife moves through an object, but the object does not split. To address this, we adopt the video-based multi-question answering pipeline, specifically to ensure the motion’s impact is checked explicitly (e.g., “After slurping, is the amount of noodles visibly reduced?”). Redundant questioning (e.g., 1. “Does the person appear to be slurping noodles?”, 2. “Is the person’s mouth in contact with the noodles?”, 3. “Is the bowl of noodles remaining after slurping?”) helps filter out false positives and accidental misinterpretations.
(5-b) Instance Preservation. Video generative models frequently generate videos with unnatural entity merging, duplication, or disappearance, particularly during large-motion sequences or object interactions. As current VLMs could not capture this clip-level or even frame-level abnormal ( i.e., Unnatural merging can appear between adjacent frames, despite both frames being normal on their own). We tune a clip-level entity abnormal detector based on Qwen2.5-VL-3B-Instruct [90]. Specifically, we use a multiple-choice question format to ask the VLM to determine whether a given clip contains any of the aforementioned anomalies (A for “yes” and B for “no”). Only the all of the clips in a video is normal will be considered as normal (score 1), otherwise 0.
In terms of training data, we first collected a set of videos that align with our prompt theme as normal samples. We also sampled several generative models (CogVideo, HunyuanVideo, Wan2.1, StepVideo), and manually annotated anomalous samples. To reduce the gap between virtual and real data, both virtual and real normal data is used for training and we further apply LoRA fine-tuning to maintain the model’s generalization ability as much as possible.
III-B Prompt Suite
By following these structured design principles, VBench-2.0 provides a compact, diverse, and reliable benchmark for evaluating video generation models across a diverse range of real-world and abstract scenarios.
III-C Human Preference Annotation
Data Preparation. There are two types of annotation formats in VBench-2.0.
Labeling Instructions. The annotation process follows VBench but incorporates refinements in interface design and evaluation methodology. Since VBench-2.0 introduces multiple dimensions with detailed multi-question evaluations and long text prompts (some exceeding 150 words), we enhance the interface for improved readability and efficiency. Instead of displaying extensive text in video titles, we sequentially list all key annotation instructions directly within the interface. This structured layout allows annotators to efficiently reference the necessary details while conducting evaluations.
Win Ratio. Given human annotations, we calculate the win ratio for each model. During pairwise comparisons, if a model’s video is preferred, it scores 1, while the other model scores 0. In case of a tie, both models score 0.5. The win ratio for each model is computed as the total score divided by the number of pairwise comparisons. We show the result in Table IV.
Quality Assurance. The annotation process maintains the rigorous quality control measures established in VBench while further refining the review criteria. We randomly sample 20% of the annotated pairs for verification, with a required success rate of 95%.
To quantify the annotation effort across all 18 evaluation dimensions, we account for the cumulative time spent on documentation drafting, initial trials, formal annotation, and re-annotation. In total, the process required 15.75 hours of individual effort and 284 hours across 18 annotators, reflecting the scale of the task and our commitment to ensuring high-quality human preference annotations. Most evaluation dimensions go through 2 rounds of trial labeling, official labeling, and post-labeling verification.